there are three forces acting on a 2kg object located at the origin. Find the values of ax and ay, the x- and y- components of the acceleration. The first force is up and to the left. It acts with -2.9N in the x direction and 3.5N in the y direction. The second force acts with 1.6N directly in the x direction. The third force acts directly in the -y direction with 2.5N of force. What are ax and ay?
Answer #1

net force in the x-direction = -2.9 + 1.6 = -1.3 N

thus, acceleration in the x- direction, ax = net force in the x-direction/mass of the object = -1.3/2 = -0.65 m/s2

net force in the x-direction = 3.5 - 2.5 = 1.0 N

thus, acceleration in the x- direction, ax = net force in the x-direction/mass of the object = 1.0/2 = 0.5 m/s2
Answer #2

ax = (-2.9 + 1.6 ) / 2 = -0.65 m /s2

ay = (3.5 - 2.5 ) /2 = 0.5 m/s2
Answer #3

to find ax right all the x component of force with ther proper sign

F1= -2.9x + 3.5y

F2 =1.6x

F3 = -2.5y ( - because it is given in the question that it is in -y direction)

Fx= max

2ax=-2.9 +1.6 =1.3

ax=1.3/2=0.65

Fy = may

2ay=3.5-2.5=1

ay = 1/2 = 0.5

so answer is

ax = 0.65

ay = 0.5